OpenStudy (anonymous):

will reward MEDAL for help! standard form for the equation of a circle? 2x^2+2y^2+8x+7=0

OpenStudy (anonymous):

\[2x ^{2}+2y ^{2}+8x+7=0\]

OpenStudy (lucaz):

you want the in form (x-a)^2+(y-b)^2=r^2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You have to create perfect squares with the numbers given. Can you try that?

OpenStudy (lucaz):

with these 2 coefficients ?

OpenStudy (anonymous):

\[2(x ^{2}+8x+16)+2y ^{2}=35\]

OpenStudy (anonymous):

You can observe that if you add +1 to each side of the equation you will have: \[2x^2+8x+2y^2+7+1=1\] which is equal to: \[2x^2+8x+2y^2+8=1\] Common factor is 2 so we have: \[2(x^2+4x+4+y^2)=1\] You can observe that \[x^2+4x+4=(x+2)^2\] Divide both sides by 2 and you will have: \[(x+2)^2+y^2=1/2\] and \[1/2=(\sqrt{1/2})^2\]. Does that help?

OpenStudy (anonymous):

ok, how do you find the radius?

OpenStudy (anonymous):

The equation is \[(x+2)^2+y^2=r^2\] where in this case, \[r=\sqrt{1/2}\] You got it?

OpenStudy (anonymous):

yes, thanks

OpenStudy (anonymous):

You are welcome! Keep it up! ;)